# here title TetraLoss: Improving the Robustness of Face Recognition against Morphing Attacks

### v1

@@ Abstract
Face recognition systems are widely deployed in high-security applications such as for biometric verification at border controls. Despite their high accuracy on pristine data, it is well-known that digital manipulations, such as face morphing, pose a security threat to face recognition systems. Malicious actors can exploit the facilities offered by the identity document issuance process to obtain identity documents containing morphed images. Thus, subjects who contributed to the creation of the morphed image can with high probability use the identity document to bypass automated face recognition systems. In recent years, no-reference (i.e., single image) and differential morphing attack detectors have been proposed to tackle this risk. These systems are typically evaluated in isolation from the face recognition system that they have to operate jointly with and do not consider the face recognition process. Contrary to most existing works, we present a novel method for adapting deep learning-based face recognition systems to be more robust against face morphing attacks. To this end, we introduce TetraLoss, a novel loss function that learns to separate morphed face images from its contributing subjects in the embedding space while still preserving high biometric verification performance. In a comprehensive evaluation, we show that the proposed method can significantly enhance the original system while also significantly outperforming other tested baseline methods.
@@ I INTRODUCTION
In this work, we propose a novel mechanism to modify the embedding space of an existing deep learning-based face recognition system and make them more robust against morphing attacks. This has the advantage that both the original face recognition embeddings and the modified embeddings can be considered, while still maintaining good biometric performance. Furthermore, such a modified face recognition system can still be combined with traditional stand-alone detection algorithms. The experimental evaluation of our proposed approach is only focused on the use case of morphing attacks. However, the method can likely be extended for other digital or physical manipulations. Figure LABEL:fig:arch_overview shows an overview of the proposed method which is referred to as TetraLoss.
@@ II-A Face Recognition Systems under Morphing Attacks
To evaluate the vulnerability to morphing attacks, different metrics have been proposed. In [37], the authors proposed the Mated Morph Presentation Match Rate (MMPMR), along with two variants in case of multiple images per subject, where a morphing attack is considered successful if all contributing subjects match against the morphed sample. In [46], the authors proposed a more restrictive measure which requires successful verification for all contributing subjects and all attempts. In [14], the authors proposed to extend these existing metrics by considering multiple face recognition systems under the motivation that a good morphing attack should potentially be able to fool multiple state-of-the-art face recognition systems. A problem with these metrics is that they only evaluate the vulnerability of the face recognition system to morphs independent of the false reject rate of the system. Hence, in [37], the authors proposed the Relative Impostor Attack Presentation Accept Rate (RIAPAR) 111previously called Relative Morph Match Rate (RMMR) which at a given security threshold considers both the attack success rate and the false non-match rate of the system. RIAPAR was recently introduced into the international standard ISO/IEC 30107-3 [19] and hence, in compliance with this, will be reported in this paper.
@@ II-B Detection of Face Morphing Attacks
Several detection methods have been proposed to increase the security of face recognition systems against morphing attacks. These can be categorized as either being Single Morphing Attack Detection (S-MAD)222also called no-reference detection methods or Differential Morphing Attack Detection (D-MAD). The latter generally achieve the most competitive performance [24, 40]. Several works have proposed such D-MAD algorithms using either handcrafted features, e.g., [3, 35, 39] or deep learning-based approaches e.g., [1, 41]. An alternative approach to detecting morphed images is face demorphing, where a live face image acquired at an eGate is used to revert part of the morphing process and reveal the legitimate document owner. This idea was first explored in [11] and has since been used in other work, e.g., in [12] where the authors improved the robustness of demorphing to face alterations.
@@ III-A Model Architecture
Inspired by the work in [2], we propose to add a shallow neural network which operates on top of embeddings from an existing face recognition system. The model consists of four fully connected (FCN) layers, batch normalization (BN) and the Leaky Rectified Linear Unit (Leaky ReLU) activation function. The input layer and the two hidden layers are followed by BN and the Leaky ReLU activation function whereas the output layer is only followed by BN. Given an embedding of size , the model performs a forward pass of the embedding and produces a new embedding of the same size (i.e., = 512 in this work). This multilayer perception (MLP) can be easily adopted to different embedding shapes introduced by different face recognition systems. It is likely that other architectures are capable of improving upon the results obtained in this work, but this is deliberately not explored in depth, as the intention is to propose a general loss and model architecture.
@@ IV-B Databases
The reason why only the datasets in [36] are considered and not other datasets, e.g., from Table I, is that they are either private or do not contain sufficient intra-class variations for realistically evaluating both the morphing vulnerability and the face recognition accuracy. Many morphing databases are derived from the Face Research London Lab (FRLL) dataset [6], e.g., [5, 34] and in the FRLL dataset, bona fide reference and probe images are considered from within the same capturing session with only minor differences in expression. This is also the case for many images in FERET, but FERET contains significantly more subjects and also has some intra-class variations (see e.g., 1(a) and 1(b)).
@@ IV-E Metrics
The face recognition performance will be evaluated according to the following metrics:
False Match Rate (FMR): The proportion of the completed biometric non-mated comparison trials that result in a false match.
False Non-Match Rate (FNMR): The proportion of the completed biometric mated comparison trials that result in a false non-match.
The vulnerability of a system is measured in terms of the Impostor Attack Presentation Accept Rate (IAPAR) which for a full-system evaluation is defined as the proportion of impostor attack presentations using the same presentation attack instrument (PAI) species that result in an accept [19]. As discussed in ISO/IEC 30107-3:2023 [19] the problem of only looking at the IAPAR at high security values, e.g., FMR=0.1% is that the convenience of the system in terms of FNMR is not considered. Hence, following the suggestion in [37], ISO/IEC 30107-3:2023 defined that the RIAPAR metrics should be used to evaluate the generalized performance of a face recognition system to morphing attacks. Given a threshold, , the RIAPAR is defined as:
@@
Specifically, we evaluate the RIAPAR by considering high security thresholds as recommended by FRONTEX [33]. FRONTEX recommends that for border security applications, a threshold corresponding to FMR=0.1% is used. Hence, we evaluate RIAPAR at thresholds corresponding to FMR=0.1%, FMR=0.01%, and FMR=0.001%.
@@ V Generalized Full-System Evaluation
As seen in Table III the RIAPAR scores obtained from the proposed method significantly outperform the original embeddings when evaluating the joint performance of the system in terms of vulnerability to morph and false reject rate. When comparing the original and proposed system without MAD, we can see that the RIAPAR of TetraLoss is better across almost every face recognition system and operating point. For instance, for FMR=0.1% the RIAPAR improves with more than 45 percentage points for both ResNet50 and ResNet100 using ArcFace, MagFace, and AdaFace. Similarly, significant gains are observed for FMR=0.01%. For FMR=0.001%, the performance gains in terms of RIAPAR using ResNet100 is between percentage points whereas for ResNet50 it is between percentage points expect for ArcFace where the original system is better. When evaluating both systems together with MAD, the proposed method still significantly outperforms the original system. Furthermore, the proposed system even, in some cases, improves when combined with MAD, especially for FMR=0.001%. The reason for this behaviour is that the used MAD system performs well on zero-effort (non-mated) attacks and hence complements the proposed system. Contrarily, the original system does not improve when adding the MAD module. The reason for this is that although the comparison scores for the morphs get lower for the fused system (less similar to the contributing subjects), the thresholds corresponding to the different FMRs also decrease and hence the system still accepts the morphed images. Since both the original face recognition system and MAD system work over the same embeddings this behavior makes sense. Future work could benchmark further baseline MAD algorithms; however, it should be noted that the MAD algorithm included in this work outperforms many existing algorithms, e.g., methods which use texture descriptors such as local binary patterns as well as landmark-based approaches and demorphing [41]. Nevertheless, the experiments in this work show that the proposed method can be used to improve the robustness of existing face recognition systems towards morphs and be used together with an existing state-of-the-art MAD algorithm.
@@
The improvements in terms of RIAPAR compared to the original system are achieved since TetraLoss is capable of pushing the embeddings of the morphed embeddings closer to the negative samples. To visualize this, consider two arbitrary embeddings, and , used in a biometric comparison. We can then calculate the squared difference vector, , between these two embeddings. Figure 4 shows a t-SNE visualization over these difference vectors when computed on 1000 randomly selected mated, non-mated and morph comparisons of the original AdaFace and modified TetraLoss embeddings using ResNet100 and ResNet50 backbones. The shown difference vectors are computed using the exact comparisons. Note that TetraLoss, in contrast to the baseline systems, seems to push the morph difference embeddings (blue points) closer to the non-mated distributions (red points), which is the desired behaviour. Ideally, the non-mated and morph embeddings should be completely overlapping. The issue with the embeddings of the original system is that, although visually there is a good separation between the three types of comparisons, we must consider the operating configuration of the system. In such a case, the operating threshold of the system would be defined according to the FMR of the system, and hence, both morphed and mated comparisons would be accepted into the system (as shown by the high RIAPAR in Table III). Obviously, more conservative decision thresholds could reduce the vulnerability towards morphing attacks. However, this would come at the cost of unacceptable high FNMR values [18]. Note also that the t-SNE spaces defined by TetraLoss (i.e., 3(b) and 3(d)) still contain morph embeddings closer to the mated embeddings. Since TetraLoss is based on the triplet loss, it is limited by the number of quadruples that must be created to achieve optimal performance. In our experiments, the FERET training database consists of few subjects, i.e., 529 subjects. Therefore, it is likely that the use of a larger training database can lead to a higher separation of the embedding spaces of mated samples and morph images.
@@ VI Feasibility Study
FMR=0.1 FMR=0.01 FMR=0.001 System Scenario IAPAR FNMR RIAPAR IAPAR FNMR RIAPAR IAPAR FNMR RIAPAR ArcFace Triplet 0.08 0.17 0.25 0.03 0.36 0.39 0.01 0.56 0.57 Triplet 0.08 0.17 0.24 0.02 0.42 0.44 0.01 0.67 0.68 Tetra 0.27 0.01 0.28 0.14 0.03 0.17 0.06 0.10 0.16 Tetra 0.28 0.00 0.29 0.14 0.02 0.17 0.05 0.11 0.16 MagFace Triplet 0.09 0.16 0.25 0.03 0.37 0.39 0.01 0.60 0.60 Triplet 0.05 0.28 0.33 0.01 0.66 0.67 0.00 0.86 0.87 Tetra 0.27 0.01 0.28 0.14 0.04 0.18 0.06 0.11 0.17 Tetra 0.24 0.01 0.25 0.09 0.07 0.15 0.02 0.25 0.27 AdaFace Triplet 0.14 0.08 0.22 0.05 0.21 0.26 0.01 0.46 0.47 Triplet 0.17 0.05 0.22 0.08 0.13 0.21 0.03 0.27 0.30 Tetra 0.27 0.01 0.28 0.15 0.04 0.18 0.05 0.12 0.18 Tetra 0.28 0.01 0.29 0.16 0.03 0.19 0.06 0.11 0.17
FMR=0.1 FMR=0.01 FMR=0.001 System Scenario IAPAR FNMR RIAPAR IAPAR FNMR RIAPAR IAPAR FNMR RIAPAR ArcFace Triplet 0.08 0.23 0.31 0.02 0.44 0.46 0.01 0.63 0.64 Triplet 0.14 0.10 0.24 0.06 0.25 0.31 0.02 0.43 0.46 Tetra 0.22 0.04 0.26 0.11 0.11 0.21 0.03 0.29 0.32 Tetra 0.25 0.03 0.27 0.13 0.08 0.21 0.06 0.18 0.24 MagFace Triplet 0.11 0.15 0.26 0.04 0.36 0.40 0.01 0.55 0.56 Triplet 0.12 0.11 0.24 0.04 0.29 0.33 0.01 0.53 0.54 Tetra 0.25 0.01 0.26 0.11 0.06 0.17 0.04 0.19 0.23 Tetra 0.26 0.01 0.28 0.14 0.05 0.19 0.05 0.15 0.20 AdaFace Triplet 0.11 0.15 0.26 0.05 0.30 0.34 0.02 0.47 0.48 Triplet 0.14 0.09 0.24 0.06 0.21 0.27 0.03 0.38 0.40 Tetra 0.24 0.02 0.26 0.12 0.07 0.20 0.06 0.16 0.22 Tetra 0.26 0.01 0.27 0.14 0.05 0.19 0.06 0.14 0.20
@@
Triplet: is calculated for each batch
@@
Tetra: is calculated for each batch
The results when training on FERET and testing on FRGCv2 for ResNet100 and ResNet50 are shown in Table IV. Note that the use of TetraLoss over the triplet loss achieves, in general, the best performance in terms of FNMR and RIAPAR, especially for high security thresholds. As expected, the IAPAR using triplet loss is in general better. However, a problem with triplet loss is that the FNMR rate becomes very high, e.g., 17% at a FMR=0.1% for ArcFace using ResNet100. This indicates the case where one out of a thousand is falsely accepted, while 17% of the mated comparisons are rejected by the system. The FNMR increases to 0.56 and 0.67 at FMR=0.001% for the Triplet and Triplet scenarios, respectively. Such a high number of false non-matches would make the system inapplicable in crowded border control situations, as it would require manual checks of the documents. Contrarily, for ArcFace using ResNet100, both TetraLoss variants have an FNMR of approximately 1% or lower at FMR=0.1%. Similar trends is observable for MagFace and AdaFace using the ResNet100 and ResNet50 backbones.

### v2

@@ Abstract
Face recognition systems are widely deployed in high-security applications such as for biometric verification at border controls. Despite their high accuracy on pristine data, it is well-known that digital manipulations, such as face morphing, pose a security threat to face recognition systems. Malicious actors can exploit the facilities offered by the identity document issuance process to obtain identity documents containing morphed images. Thus, subjects who contributed to the creation of the morphed image can with high probability use the identity document to bypass automated face recognition systems. In recent years, no-reference (i.e., single image) and differential morphing attack detectors have been proposed to tackle this risk. These systems are typically evaluated in isolation from the face recognition system that they have to operate jointly with and do not consider the face recognition process. Contrary to most existing works, we present a novel method for adapting deep learning-based face recognition systems to be more robust against face morphing attacks. To this end, we introduce TetraLoss, a novel loss function that learns to separate morphed face images from its contributing subjects in the embedding space while still achieving high biometric verification performance. In a comprehensive evaluation, we show that the proposed method can significantly enhance the original system while also significantly outperforming other tested baseline methods.
@@ I INTRODUCTION
In this work, we propose a novel mechanism to modify the embedding space of an existing deep learning-based face recognition system and make them more robust against morphing attacks. This has the advantage that both the original face recognition embeddings and the modified embeddings can be considered, while still maintaining good biometric performance. Furthermore, such a modified face recognition system can still be combined with traditional stand-alone detection algorithms. The experimental evaluation of our proposed approach is only focused on the use case of morphing attacks. However, the method can likely be extended for other digital or physical manipulations. Figure LABEL:fig:arch_overview shows an overview of the proposed method which is referred to as TetraLoss111https://github.com/dasec/TetraLoss.
@@ II-A Face Recognition Systems under Morphing Attacks
To evaluate the vulnerability to morphing attacks, different metrics have been proposed. In [39], the authors proposed the Mated Morph Presentation Match Rate (MMPMR), along with two variants in case of multiple images per subject, where a morphing attack is considered successful if all contributing subjects match against the morphed sample. In [48], the authors proposed a more restrictive measure which requires successful verification for all contributing subjects and all attempts. In [15], the authors proposed to extend these existing metrics by considering multiple face recognition systems under the motivation that a good morphing attack should potentially be able to fool multiple state-of-the-art face recognition systems. A problem with these metrics is that they only evaluate the vulnerability of the face recognition system to morphs independent of the false reject rate of the system. Hence, in [39], the authors proposed the Relative Impostor Attack Presentation Accept Rate (RIAPAR) 222Previously called Relative Morph Match Rate (RMMR) which at a given security threshold considers both the attack success rate and the false non-match rate of the system. RIAPAR was recently introduced into the international standard ISO/IEC 30107-3 [21] and hence, in compliance with this, will be reported in this paper.
@@ II-B Detection of Face Morphing Attacks
Several detection methods have been proposed to increase the security of face recognition systems against morphing attacks. These can be categorized as either being Single Morphing Attack Detection (S-MAD)333Also called no-reference detection methods or Differential Morphing Attack Detection (D-MAD). The latter generally achieve the most competitive performance [26, 42]. Several works have proposed such D-MAD algorithms using either handcrafted features, e.g., [3, 37, 41] or deep learning-based approaches e.g., [1, 43]. An alternative approach to detecting morphed images is face demorphing, where a live face image acquired at an eGate is used to revert part of the morphing process and reveal the legitimate document owner. This idea was first explored in [12] and has since been used in other work, e.g., in [13] where the authors improved the robustness of demorphing to face alterations.
@@ III-A Model Architecture
Inspired by the work in [2], we propose to add a shallow neural network which operates on top of embeddings from an existing face recognition system. The model consists of four fully connected (FCN) layers, batch normalization (BN) and the Leaky Rectified Linear Unit (Leaky ReLU) activation function. The input layer and the two hidden layers are followed by BN and the Leaky ReLU activation function whereas the output layer is only followed by BN. Given an embedding of size , the model performs a forward pass of the embedding and produces a new embedding of the same size (i.e., = 512 in this work). This multilayer perception (MLP) can be easily adopted to different embedding shapes introduced by different face recognition systems. It is likely that other architectures are capable of improving upon the results obtained in this work, but this is deliberately not explored in depth, as the intention is to propose a general loss and model architecture. Applying a MLP on top of embeddings extracted from an existing face recognition system can be thought of as a type of adapter where the embeddings are adapted for a new task.
@@ IV-B Databases
The reason why only the datasets in [38] are considered and not other datasets, e.g., from Table I, is that they are either private or do not contain sufficient intra-class variations for realistically evaluating both the morphing vulnerability and the face recognition accuracy. Many morphing databases are derived from the Face Research London Lab (FRLL) dataset [7], e.g., [5, 36] and in the FRLL dataset, bona fide reference and probe images are considered from within the same capturing session with only minor differences in expression. This is also the case for many images in FERET, but FERET contains significantly more subjects and also has some intra-class variations (see e.g., 2(a) and 2(b)).
@@ IV-E Metrics
The face recognition performance will be evaluated according to the False Match Rate (FMR) and False Non-Match Rate (FNMR) as further described in [20]. Furthermore, the vulnerability of a system is measured in terms of the Impostor Attack Presentation Accept Rate (IAPAR) which for a full-system evaluation is defined as the proportion of impostor attack presentations using the same presentation attack instrument (PAI) species that result in an accept [21]. As discussed in ISO/IEC 30107-3:2023 [21] the problem of only looking at the IAPAR at high security values, e.g., FMR=0.1% is that the convenience of the system in terms of FNMR is not considered. Hence, following the suggestion in [39], ISO/IEC 30107-3:2023 defined that the RIAPAR metrics should be used to evaluate the generalized performance of a face recognition system to morphing attacks. Given a threshold, , the RIAPAR is defined as:
@@
Specifically, we evaluate the RIAPAR by considering high security thresholds as recommended by FRONTEX [35]. FRONTEX recommends that for border security applications, a threshold corresponding to FMR=0.1% is used. Hence, we evaluate RIAPAR at thresholds corresponding to FMR=0.1%, FMR=0.01%, and FMR=0.001%. During the experiments, the thresholds corresponding to the above FMR values are calculated from scores obtained by doing more than two million bona fide non-mated comparisons on the FRGCv2 database.
@@ V Generalized Full-System Evaluation
As seen in Table III the RIAPAR scores obtained from the proposed method significantly outperform the original embeddings when evaluating the joint performance of the system in terms of vulnerability to morph and false reject rate. When comparing the original and proposed system without MAD, we can see that the RIAPAR of TetraLoss is better across almost every face recognition system and operating point. For instance, for FMR=0.1% the RIAPAR improves with more than 45 percentage points for both ResNet50 and ResNet100 using ArcFace, MagFace, and AdaFace. Similarly, significant gains are observed for FMR=0.01%. For FMR=0.001%, the performance gains in terms of RIAPAR using ResNet100 is between percentage points whereas for ResNet50 it is between percentage points expect for ArcFace where the original system is better. When evaluating both systems together with MAD, the proposed method still significantly outperforms the original system. Furthermore, the proposed system even, in some cases, improves when combined with MAD, especially for FMR=0.001%. The reason for this behaviour is that the used MAD system performs well on zero-effort (non-mated) attacks and hence complements the proposed system. Contrarily, the original system does not improve when adding the MAD module. The reason for this is that although the comparison scores for the morphs get lower for the fused system (less similar to the contributing subjects), the thresholds corresponding to the different FMRs also decrease and hence the system still accepts the morphed images. Since both the original face recognition system and MAD system work over the same embeddings this behavior makes sense. Future work could benchmark further baseline MAD algorithms; however, it should be noted that the MAD algorithm included in this work outperforms many existing algorithms, e.g., methods which use texture descriptors such as local binary patterns as well as landmark-based approaches and demorphing [43]. It should also be noted that for all systems and backbones at the three selected operating points, the original system retains a FNMR close to 0% whereas the FNMR of the TetraLoss system gradually increases and hence degrades performance on bona fide data (see Table IV). In general, the FNMR of the original system will also increase if the security threshold is chosen based on even lower FMR values than evaluated herein (as shown in [19] on similar data). Future work could focus on more challenging databases with more intra-class variation and using a more significant number of morphed images. Nevertheless, the experiments in this work indicate that, at relevant operating points, the proposed method can improve the robustness of existing face recognition systems towards morphs and be used with an existing state-of-the-art MAD algorithm.
@@
The improvements in terms of RIAPAR compared to the original system are achieved since TetraLoss is capable of pushing the embeddings of the morphed embeddings closer to the negative samples. To visualize this, consider two arbitrary embeddings, and , used in a biometric comparison. We can then calculate the squared difference vector, , between these two embeddings. Figure 4 shows a t-SNE visualization over these difference vectors when computed on 1000 randomly selected mated, non-mated and morph comparisons of the original AdaFace and modified TetraLoss embeddings using ResNet100 and ResNet50 backbones. The shown difference vectors are computed using the exact comparisons. Note that TetraLoss, in contrast to the baseline systems, seems to push the morph difference embeddings (blue points) closer to the non-mated distributions (red points), which is the desired behaviour. Ideally, the non-mated and morph embeddings should be completely overlapping. The issue with the embeddings of the original system is that, although visually there is a good separation between the three types of comparisons, we must consider the operating configuration of the system. In such a case, the operating threshold of the system would be defined according to the FMR of the system, and hence, both morphed and mated comparisons would be accepted into the system (as shown by the high RIAPAR in Table III). Obviously, more conservative decision thresholds could reduce the vulnerability towards morphing attacks. However, this would come at the cost of unacceptable high FNMR values [19]. Note also that the t-SNE spaces defined by TetraLoss (i.e., 4(b) and 4(d)) still contain morph embeddings closer to the mated embeddings. Since TetraLoss is based on the triplet loss, it is limited by the number of quadruples that must be created to achieve optimal performance. In our experiments, the FERET training database consists of few subjects, i.e., 529 subjects. Therefore, it is likely that the use of a larger training database can lead to a higher separation of the embedding spaces of mated samples and morph images.
@@ VI Feasibility Study
FMR=0.1 FMR=0.01 FMR=0.001 System Scenario IAPAR FNMR RIAPAR IAPAR FNMR RIAPAR IAPAR FNMR RIAPAR ArcFace Triplet 0.08 0.17 0.25 0.03 0.36 0.39 0.01 0.56 0.57 Triplet2 0.08 0.17 0.24 0.02 0.42 0.44 0.01 0.67 0.68 Tetra 0.27 0.01 0.28 0.14 0.03 0.17 0.06 0.10 0.16 Tetra2 0.28 0.00 0.29 0.14 0.02 0.17 0.05 0.11 0.16 MagFace Triplet 0.09 0.16 0.25 0.03 0.37 0.39 0.01 0.60 0.60 Triplet2 0.05 0.28 0.33 0.01 0.66 0.67 0.00 0.86 0.87 Tetra 0.27 0.01 0.28 0.14 0.04 0.18 0.06 0.11 0.17 Tetra2 0.24 0.01 0.25 0.09 0.07 0.15 0.02 0.25 0.27 AdaFace Triplet 0.14 0.08 0.22 0.05 0.21 0.26 0.01 0.46 0.47 Triplet2 0.17 0.05 0.22 0.08 0.13 0.21 0.03 0.27 0.30 Tetra 0.27 0.01 0.28 0.15 0.04 0.18 0.05 0.12 0.18 Tetra2 0.28 0.01 0.29 0.16 0.03 0.19 0.06 0.11 0.17
FMR=0.1 FMR=0.01 FMR=0.001 System Scenario IAPAR FNMR RIAPAR IAPAR FNMR RIAPAR IAPAR FNMR RIAPAR ArcFace Triplet 0.08 0.23 0.31 0.02 0.44 0.46 0.01 0.63 0.64 Triplet2 0.14 0.10 0.24 0.06 0.25 0.31 0.02 0.43 0.46 Tetra 0.22 0.04 0.26 0.11 0.11 0.21 0.03 0.29 0.32 Tetra2 0.25 0.03 0.27 0.13 0.08 0.21 0.06 0.18 0.24 MagFace Triplet 0.11 0.15 0.26 0.04 0.36 0.40 0.01 0.55 0.56 Triplet2 0.12 0.11 0.24 0.04 0.29 0.33 0.01 0.53 0.54 Tetra 0.25 0.01 0.26 0.11 0.06 0.17 0.04 0.19 0.23 Tetra2 0.26 0.01 0.28 0.14 0.05 0.19 0.05 0.15 0.20 AdaFace Triplet 0.11 0.15 0.26 0.05 0.30 0.34 0.02 0.47 0.48 Triplet2 0.14 0.09 0.24 0.06 0.21 0.27 0.03 0.38 0.40 Tetra 0.24 0.02 0.26 0.12 0.07 0.20 0.06 0.16 0.22 Tetra2 0.26 0.01 0.27 0.14 0.05 0.19 0.06 0.14 0.20
@@
Triplet2: is calculated for each batch
@@
Tetra2: is calculated for each batch
The results when training on FERET and testing on FRGCv2 for ResNet100 and ResNet50 are shown in Table IV. Note that the use of TetraLoss over the triplet loss achieves, in general, the best performance in terms of FNMR and RIAPAR, especially for high security thresholds. As expected, the IAPAR using triplet loss is in general better. However, a problem with triplet loss is that the FNMR rate becomes very high, e.g., 17% at a FMR=0.1% for ArcFace using ResNet100. This indicates the case where one out of a thousand is falsely accepted, while 17% of the mated comparisons are rejected by the system. The FNMR increases to 0.56 and 0.67 at FMR=0.001% for the Triplet and Triplet2 scenarios, respectively. Such a high number of false non-matches would make the system inapplicable in crowded border control situations, as it would require manual checks of the documents. Contrarily, for ArcFace using ResNet100, both TetraLoss variants have an FNMR of approximately 1% or lower at FMR=0.1%. Similar trends is observable for MagFace and AdaFace using the ResNet100 and ResNet50 backbones.